OpenStudy (anonymous):

7(2x+45) = 693-45x Find x I give medals!

OpenStudy (igreen):

Simplify it first. \(7(2x + 45) = 693 - 45x\) Distribute 7 into the parenthesis: \(14x + 315 = 693 - 45x\) Can you finish that? @QuestionSheep

OpenStudy (anonymous):

From my screen it says "[Math Proscessing Error]"

OpenStudy (igreen):

Oh..okay. Well: here's what I wrote: Simplify it first. 7(2x + 45) = 693 - 45x Distribute 7 into the parenthesis: 14x + 315 = 693 - 45x Can you finish that?